The charge of a body is +1.6 × 10^20C.Which of the following can be drawn from the given statement.1-The body has 10^39 electrons2-The body has 10^39 protons3-The body has lost 10^39 electrons4-The body has gain 10^39 electrons.

Grade:8

1 Answers

Naman Jain
6 years ago
the body has 10^39 electrons.
n = q/e      [ n = no. of electrons, q = charge, e = charge of electron ]
n = 1.6×10^20
      1.6×10^-19
  n = 10^39